# On-chip microwave coherent source with in-situ control of the photon number distribution

### v2

@@ II.2 Cavity control through tunable coupling
Figure 3: Coherent photon generation through the tunable coupler scheme. (a) Flux timing diagram illustrating the photon population protocol of a single reservoir with the proposed scheme. At the initial time $\mathrm{t_{\Gamma}}$ the counter-relaxation is switched on. Afterwards the flux $\mathrm{\Phi_{c}}$ on the tunable coupler is chosen according to the wanted steady-state distribution of photons in the reservoir. After the time $t_{ss}$ where the system has reached its steady state, the photon source is set off resonance from the reservoir to avoid further interactions between the two systems. (b) Steady-state photon distributions in the target reservoir under varying flux $\Phi_{c}$ values threading the coupler. These simulation results are obtained for a device engineered according to the parameters listed at the end of paragraph \@slowromancapii@B. Each color represents the distribution of photons in the Fock basis corresponding to specific values of the flux $\mathrm{\Phi_{c}}$. (c) Steady-state reservoir normalized Wigner functions for different values of the flux $\mathrm{\Phi_{c}}$. The radius of the Wigner function in the phase space is proportional to the average of the corresponding photon number distribution
injected into the cavity can span in the entire range from 0 to $\mathrm{\Gamma/2\kappa}$. The dark blue area, conversely, contains the working points where there is no degree of control via the external flux in the considered range of effective coupling. Even when the system is well designed to operate in an appropriate working point, it must also possess robustness against fabrication errors, which can impact the position of the working point on the map, and against flux noise in the system, which can negatively affect the control over the steady-state photons distribution. To prevent the influence of faulty fabrication on the system feasibility, the coupling can be engineered to work within a broader range. This expansion of the coupling range, however, comes at the expense of an increased sensitivity of the tunable coupler to external flux. In a practical realization of the system, a trade-off must be found between these two features. To decrease the sensitivity to external flux, we can employ several strategies that allow to define the range of coupling of the system precisely. Differently from the typical tunable coupler scheme in which the system is engineered to have an off-point value where the coupling is canceled [32], here we just aim to modulate it within a specific range. One approach involves designing the tunable coupler with a resonance frequency lower than that of the photon source and the reservoir, resulting in an effective coupling strength $\mathrm{g_{eff}\propto 1/(\omega_{s}-\omega_{c})}$. Additionally, we can use an asymmetric SQUID as the tunable element of the tunable coupler, which exhibits a more gradual response to external fluxes [54]. These adjustments provide two significant advantages: (1) the coupler never approaches a frequency close to those of the photon source and the reservoir for any threaded flux, ensuring that the dispersive condition is consistently met; (2) the asymmetry allows for non-zero minima in the coupler frequency, narrowing the range of effective coupling to well-defined maxima and minima. Alternative solutions rely on employing a pulsed approach [55]. To demonstrate the feasibility of the proposed scheme, we consider the following system: a flux-tunable photon source with a counter-relaxation rate $\Gamma=10\,\mathrm{MHz}$, self-capacity of $C_{s}=100\,\mathrm{fF}$, and a transition frequency $\omega_{s}=5\,\mathrm{GHz}$ that is resonant with a reservoir having a resonance frequency $\omega_{r}=\omega_{s}$, self-capacity of $C_{r}=400\,\mathrm{fF}$, and loss rate $\kappa=0.1\,\mathrm{MHz}$. The two are connected through a tunable coupler corresponding to an asymmetric flux-tunable qubit with Josephson energies $E_{J1}^{c}=31.4\,\mathrm{GHz}$, $E_{J2}^{c}=37.7\,\mathrm{GHz}$, and capacitive energy $E_{C}^{c}=70\,\mathrm{MHz}$. The coupling capacities between the photon source and the reservoir with the tunable coupler are $C_{sc}=C_{rc}=1\,\mathrm{fF}$. A design with these parameters is perfectly feasible for state of the art nanofabrication process and the value of $\Gamma$ is near the one used in other works [50, 30]. We analyze the system dynamics for various values of the flux $\mathrm{\Phi_{c}}$ threading the coupler. By activating the pumping process and exploring the system across different flux values, we obtain the results shown in Fig. 3(b). Different values of the flux threading the coupler yield distinct values of the effective coupling $\mathrm{g_{eff}}$, which ultimately lead to different photon distributions within the reservoir at the steady state. The Wigner function defining the steady state of the reservoir for different values of the effective coupling are shown in Fig. 3(c). We observe the typical ”donut” shape that characterizes coherent photon sources, where the radius of the Wigner function is proportional to the average of the steady-state photon distribution. After the reservoir has reached its steady state, the pumping and the flux threading the tunable coupler can be switched off. At this stage, to avoid further interactions between the photon source and the reservoir we can change the transition frequency of the photon source through the external flux $\mathrm{\Phi_{s}}$.
II.3 Multiple resonator generalization
@@
Another advantage offered by the proposed scheme is the ability to address multiple cavities by exploiting the tunability of the photon source. Under the same assumptions as the single-reservoir case, we can extend the analysis to an $\mathrm{N_{r}}$-reservoir system. The system is described by a Hamiltonian that includes additional terms related to the self-energy of different reservoirs, the coupling of each reservoir to the tunable coupler, and the second nearest neighbor cross-couplings between each resonator. In this discussion, we consider the simple case of a two-reservoir setup as a starting point, where the case of $\mathrm{N_{r}}$ reservoirs will be a direct generalization (calculations of the general $\mathrm{N_{r}}$ case are provided in the Appendix). We investigate the ability to individually address each reservoir through the external fluxes, threading the tunable coupler and assuming that the photon source can be tuned on resonance with both resonators separately. Assuming that both the coupler and the photon source behave as perfect two-level systems and neglecting second-order interactions, the Hamiltonian is written as:
$\mathrm{\hat{H_{N}}=\hat{H}_{s}+\hat{H}_{c}+\hat{H}_{t}+\hat{H}_{i}+\sum_{j=s,r_{1},r_{2}}\hat{H}_{jc}+\hat{H}_{ti}}\,.$ (6)
@@
$\mathrm{\frac{\hat{H}}{\hbar}}$ $=\mathrm{\frac{\tilde{\omega_{s}}\sigma_{s}^{z}}{2}+\tilde{\omega_{r_{1}}}a_{r_{1}}+\tilde{\omega_{r_{2}}}a_{r_{2}}^{+}a_{r_{2}}}$ (8)
$\mathrm{\phantom{=}+\sum_{j=r_{1},r_{2}}\tilde{g}_{jc}(\sigma_{s}^{+}a_{j}+\sigma_{s}^{-}a_{j}^{+})}$
$\mathrm{\phantom{=}+\tilde{g}_{r_{1}r_{2}}(a_{r_{1}}a^{+}_{r_{2}}+a_{r_{1}}^{+}a_{r_{2}})}\,.$
In this expression, the renormalizations of various frequency terms are identical to those observed in the single-reservoir case, except for the new cross-coupling term $\mathrm{\tilde{g}_{r_{1}r_{2}}}$ between the two reservoirs. The second-order coupling term $\mathrm{g_{r_{1}r_{2}}}$ is summed with another term arising from the coupling between the reservoirs via the tunable coupler, leading to the following expression of the effective cross-coupling strength:
@@
It is, therefore, crucial to understand how much detuning is required between the different reservoirs at the fan-out of the system to avoid resonant interactions that lead to cross-population between the cavities. We consider a setup similar to the single-reservoir scheme analyzed in the previous section, with the addition of a second cavity at a frequency $\mathrm{\omega_{r_{2}}=\omega_{r_{1}}+\Delta}$ and linked to the tunable coupler with the same capacitive coupling used for the target cavity. We observe the system behavior for different values of $\mathrm{\Delta}$ while sweeping both the photon source and the coupler frequencies, simulating the effect of the external fluxes $\mathrm{\Phi_{s}}$ and $\Phi_{c}$ individually threading the two elements, respectively. Assuming that both reservoirs have the same linewidth as the reservoir considered in the previous single reservoir case $\mathrm{\kappa_{1}=\kappa_{2}=0.1MHz}$. To avoid cross-population between the two reservoirs the detuning $\mathrm{\Delta=\omega_{r_{1}}-\omega_{r_{2}}}$ must be much higher than the spectral linewidth of the two reservoirs $\mathrm{(\Delta\gg\kappa)}$. In this case, we chose a detuning of $\mathrm{\Delta/2\pi\geq 5\,MHz}$, that allows to individually address and control the steady-state population of each resonator with the proposed scheme. In Fig. 4(a), we show the behavior of the system, where by setting the photon source on resonance with one of the cavities, it is possible to control their population without populating also the other resonator. For small detuning, the effective coupling between the source and the cavity has approximately the same dependence on the flux, even for a fixed value of the capacitive coupling between the cavity and the tunable coupler. When engineering systems are characterized by high detuning between the different reservoirs on the output, the respective capacitive links to the tunable coupler must be adequately designed. To optimize the device, the couplings must be chosen to preserve the shape of $\mathrm{g_{eff}(\Phi_{c})}$, because the effective coupling depends on the resonance frequency of the target reservoir $\mathrm{\omega_{t}}$. We thus need to take into account the variation of this frequency by changing the coupling between the central coupler and the cavity under consideration. We use the same setup employed before, taking a reference cavity, with a resonance frequency $\mathrm{\omega_{r_{1}}}$ and a capacity $\mathrm{C_{r_{1}c}}$ linking the cavity to the central coupler. We consider a second cavity with a detuned resonance frequency of $\mathrm{\omega_{r_{2}}=\omega_{r_{1}}+\Delta}$ and with a coupling capacitor $\mathrm{C_{r_{2}c}=\alpha C_{r_{1}c}}$, as depicted in Fig. 1(a). We assume that the self-capacity of the cavity does not change with its resonance frequency. If there is a change it will result in an additional factor on the expression of the coupling strength that can be absorbed by the factor $\alpha$, allowing the same type of analysis. Through a sweep on $\mathrm{\Delta}$ and $\mathrm{\alpha}$, we study how the shape of $\mathrm{g_{eff}(\Phi_{c})}$ changes by analyzing the Hausdorff distance relative to the reference curve obtained for the reference cavity. The results of this analysis are shown in Fig. 4(b), where we can clearly distinguish a blue region corresponding to the optimal values of the coupling capacity. For small values of the detuning $\mathrm{\Delta}$, the ideal coupling correction $\mathrm{\alpha}$ is near unity and, in general, follows the shape $\mathrm{\alpha=\sqrt{\omega_{r}/(\omega_{r}+\Delta)}}$.
Figure 4: Multiple reservoirs case photon population. (a) Steady-state photon number in different cavities with resonance frequencies $\mathrm{\omega_{r_{1}}/2\pi=5GHz}$ and $\mathrm{\omega_{r_{2}}/2\pi=5.005GHz}$. The red and blue points on the two plots represent the average number of injected photons in each cavity for specific choices of flux $\mathrm{\Phi_{c}}$ and frequency $\mathrm{\omega_{s}/2\pi}$ of the photon source. The absence of colored points in correspondence to the frequency of the other resonator in each colorplot shows the absence of cross-population phenomena. The rest of the parameters are the same as the one considered in Fig. 3. (b) The Hausdorff distance between the effective coupling strength curve $\mathrm{g_{eff}(\Phi_{c})}$ of the reference cavity and the second cavity detuned by $\mathrm{\Delta}$ and coupled with a coupling capacity $\mathrm{\alpha}$ times that of the reference cavity. The dark blue zone indicates a range of coupling capacities that maintain consistent effective coupling dependencies while altering the resonance frequency of the reservoir relative to the reference cavity.
III Conclusions
@@
We have proposed a simple and versatile scheme for an on-demand coherent photon source on chip. The tunability inherent to this scheme enables precise control over the average of the steady-state photon distribution generated in a target reservoir while also facilitating the addressing of different cavities capacitively linked to the coupler, each characterized by distinct resonance frequencies. Initially, we consider the case of a single resonator, where after identifying a set of parameters for a practical design we have analyzed the system performance for various fluxes applied across the coupler, demonstrating the capability of the proposed setup to control the photon distribution generated in a target cavity. Subsequently, we extend our analysis to a setup comprising an additional reservoir, showcasing the capability to individually address different reservoirs at different frequencies by adequately designing the detunings and the couplings in the system. The device acts as a tunable and finely controllable cryogenic micromaser, which can be an important tool in many quantum architectures based on bosonic quantum systems, [34, 35, 36], like continuous variable quantum computing [37], quantum memories [56] and quantum metrology [57]. Moreover, the demultiplexing capability of the proposed scheme offers the advantage of decoupling multiple reservoirs from direct connections to room-temperature devices, leading to significant enhancements in terms of scalability. The scheme is also well-suited for 3D integrated systems, further expanding its domain of potential applications.
IV Acknowledgments
The authors thank F. Nori and A. Porzio for the fruitful discussion and would also like to thank G. Di Bello and M. Vizzuso for their help in the work development. The work was supported by the project “On-chip signal generation for superconducting quantum processors (SFQ4QPU)”, in the frame of Eurostars CoD 15 Call 2021, by the project “Superconducting quantum-classical linked computing systems (SuperLink)”, in the frame of QuantERA2 ERANET COFUND in Quantum Technologies, and by the Project PRIN 2022-Advanced Control and Readout of scalable Superconducting NISQ Architectures (SuperNISQ)-CUP E53D23001910006. The research activities were also supported by the PNRR MUR project PE0000023-NQSTI and the PNRR MUR project CN_00000013—ICSC.
V Data availability
VI Code availability
VII Appendix
VII.1 Hamiltonian

### v3

@@ II.3 Control through tunable coupling
Figure 4: Coherent photon generation through the tunable coupler scheme. (a) Flux timing diagram illustrating the photon population protocol of a single reservoir with the proposed scheme. The flux $\Phi_{s}$ sets the emission frequency on resonance with the target reservoir. At the initial time $\mathrm{t_{\Gamma}}$ the counter-relaxation is switched on. Afterwards the flux $\mathrm{\Phi_{c}}$ on the tunable coupler is chosen according to the wanted steady-state distribution of photons in the reservoir. After the time $t_{ss}$ where the system has reached its steady state, the photon source is set off resonance from the reservoir by changing $\Phi_{s}$ to avoid further interactions between the two systems. (b) Steady-state photon distributions in the target reservoir under varying flux $\Phi_{c}$ values threading the coupler. These simulation results are obtained by simulating the steady-state dynamics in Qutip with the Hamiltonian in Eq. 6 for a device engineered according to the parameters listed at the end of Sec. Control through tunable coupling. Each color represents the distribution of photons in the Fock basis corresponding to specific values of the flux $\mathrm{\Phi_{c}}$. The dimension $\mathrm{N}$ of the Hilbert subspace associated with the reservoir in the simulation is $\mathrm{N=100}$. (c) Corresponding steady-state reservoir normalized Wigner functions for different values of the flux $\mathrm{\Phi_{c}}$. The radius of the Wigner function in the phase space is proportional to the average of the corresponding photon number distribution as in typical IQ representations of spontaneous emission from a photon source [33, 37, 21].
Following these guidelines, the effective coupling can be engineered to have the shape shown in Fig. 4(b). In this configuration, the system passes from a masing regime $(\lambda=0.1)$ to a suppressed masing regime $(\lambda=2)$ at the two extrema of the flux. Notably, at these extremal points, the sensitivity to flux is suppressed $(\partial g_{eff}/\partial\Phi\approx 0)$, making the system particularly insensitive to flux noise in these operating regimes. The system can also be operated exclusively at the two extremal points of the flux range, where the photon population can be controlled by adjusting the duration of the flux pulse applied to the tunable coupler, which switches between the points of maximum and minimum effective coupling. Alternative approaches involve the use of ferromagnetic elements, which enable flux tunability without the application of static magnetic fields commonly identified as a primary source of $1/f$ flux noise [73]. The design of the tunable coupler must also ensure that its interaction with the other elements always satisfies the dispersive approximation within the operational flux range of the system. To guarantee the general validity of this condition, the tunable coupler and the capacitive coupling to the other components must be engineered such that the number of injected photons $\mathrm{n_{ph}}$ never reach values close to the critical number of photons $\mathrm{n_{cr}=\left(\Delta_{rc}/2g_{rc}\right)^{2}}$ [63]. The critical number of photons depends on the detuning $\mathrm{\Delta_{rc}}$ between the reservoir and the coupler, and the coupling strength, $\mathrm{g_{rc}}$, between them. In scenarios where the tunable coupler is designed to always have a frequency below the resonance frequencies of the other elements, $\mathrm{n_{cr}}$ reaches its minimum at zero flux $\mathrm{\left(\Phi_{c}=0\right)}$, where the detuning is minimized. Notably, the zero flux point also corresponds to the regime of maximum photon emission, making it the point where the dispersive approximation is most likely to break.
The maximum number of photons $\mathrm{n_{ph}^{max}=\Gamma/2\kappa}$ is defined by the working point $\mathrm{\left(\Gamma,\kappa\right)}$. Therefore, the dispersive approximation holds if the ratio $\mathrm{n_{ph}^{max}/n_{cr}=2\Gamma g_{rc}^{2}\left(0\right)/\kappa\Delta_{rc}^{2}\left(0\right)\ll 1}$ in any flux point. Moreover, since by definition $n_{cr}/N^{*}=\left(\Delta_{rc}/\Delta_{min}\right)^{2}\left(g_{sr}/2g_{rc}\right)^{2}<1$, this assumption also ensures that the splitting between dressed states can be fairly neglected for any flux $\mathrm{\Phi_{c}}$ threading the tunable coupler. Given the typical weak anharmonicity $\mathrm{\alpha_{c}}$ of the transmon qubits used as tunable coupler [74, 46], it is important to account for the effects of higher-energy states on the virtual interaction between the reservoir and the source within the scheme. The two eigenstates |$\mathrm{\ket{e,0,n_{ph}}}$ and |$\mathrm{\ket{g,0,n_{ph}+1}}$, which govern photon transport from the source to the reservoir, can couple to higher-order states such as |$\mathrm{\ket{g,2,n_{ph}-1}}$, originating from the higher excited states of the tunable coupler. Assuming the photon source and reservoir are on resonance $\mathrm{\left(\omega=\omega_{s}=\omega_{r}\right)}$, the energy separation from the second excited state of the tunable coupler is approximately $\mathrm{\hbar\delta_{\ket{g,2,n_{ph}-1}}\approx 2\hbar\left(\omega-\omega_{c}\right)-\hbar\alpha_{c}}$, where we neglect the dependence on the photon number $\mathrm{n_{ph}}$ embedded in the dressed state splitting of the eigenestates. This expression suggests that a potential mitigation strategy is to design tunable couplers with higher anharmonicity.
When engineering the anharmonicity of the tunable coupler, it is also important to consider the self-Kerr effect induced by the coupler on the reservoir dynamics. The nonlinearity of the tunable coupler introduces a term $\mathrm{\hat{H}_{K}=K_{r}\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}}$ in the reservoir Hamiltonian, where $\mathrm{K_{r}}$ represents the self-Kerr of the reservoir. In this scheme, the only contribution to $\mathrm{K_{r}}$ arises from the interaction with the tunable coupler and is defined by $\mathrm{K_{r}=\chi_{rc}^{2}/4\alpha_{c}}$ [75, 76], where $\mathrm{\chi_{rc}=-g_{rc}^{2}|\alpha_{c}|/\Delta_{rc}\left(\Delta_{rc}-|\alpha_{c}|\right)}$ is the dispersive shift associated to the virtual interaction between the reservoir and the coupler [63]. To ensure that the Kerr term does not impact the time evolution of the generated state in the reservoir, the ratio $\mathrm{K_{r}/\kappa}$ must satisfy $\mathrm{K_{r}/\kappa\ll 1}$ [75]. Under this condition, the Kerr evolution is sufficiently slow to be negligible compared to the reservoir state collapse, allowing it to be effectively treated as harmonic. The implementation of the scheme for situations where Kerr dynamics dominate over the loss rate is under investigation.
To demonstrate the feasibility of the proposed scheme, we consider the dynamics of the following example scheme: a flux-tunable photon source with a counter-relaxation rate $\Gamma=50\,\mathrm{MHz}$, self-capacity of $C_{s}=100\,\mathrm{fF}$, and a transition frequency $\omega_{s}/2\pi=7\,\mathrm{GHz}$ that is resonant with a reservoir having a resonance frequency $\omega_{r}=\omega_{s}$, self-capacity of $C_{r}=500\,\mathrm{fF}$, and loss rate $\kappa/2\pi=0.5\,\mathrm{MHz}$. The two are connected through a tunable coupler corresponding to an asymmetric flux-tunable qubit with Josephson energies $E_{J1}^{c}/h=10.02\,\mathrm{GHz}$, $E_{J2}^{c}/h=8.532\,\mathrm{GHz}$, and charging energy $E_{C}^{c}/h=200\,\mathrm{MHz}$, resulting in a ratio $\mathrm{E_{j_{2}}^{c}/E_{C}^{c}\approx 43}$. The coupling capacities between the photon source and the reservoir with the tunable coupler are $C_{sc}=C_{rc}=3.5\,\mathrm{fF}$. A design with these parameters is perfectly feasible for the state-of-the-art nanofabrication process [14, 46, 54] and the value of $\mathrm{\Gamma}$ is near the one used in other works [59, 62].
A viable candidate that can be used as the photon source in this scheme is an aluminum voltage-biased charge qubit [32]. Designing the chip with a Josephson junction of energy $E_{j}^{s}/h\sim 14.66GHz$ and a shunt resistance on the electrode of $R_{b}=5\,M\Omega$ we achieve the target $\mathrm{\Gamma/2\pi\sim 50\,MHz}$. With the chosen parameters, the ratio $\mathrm{n_{ph}^{max}/n_{cr}\left(\Phi_{c}=0\right)=0.15}$ which ensures that the system is well described by the Hamiltonian in Eq. 6 in any flux point. The maximum value of the self-Kerr induced by the tunable coupler on the reservoir appears at zero flux point where $\mathrm{|K_{r}|\sim 0.0365\,kHZ\ll\kappa}$. The small value of the self-Kerr $\mathrm{K_{r}}$ compared to the loss rate $\mathrm{\kappa}$ guarantees that Kerr evolution is negligible over the lifetime of the generated state in the reservoir of the proposed scheme. We analyze the system dynamics for various values of the flux $\mathrm{\Phi_{c}}$ threading the coupler. By activating the pumping process and exploring the system across different flux values, we obtain the results shown in Fig. 4(b). Different values of the flux threading the coupler yield distinct values of the effective coupling $\mathrm{g_{eff}}$, which ultimately lead to different photon distributions within the reservoir at the steady state. Assuming typical flux fluctuations on the order of $10\,\mu\Phi_{0}$ [77, 46], the resulting variation in photon number is negligible, with deviations of the order of $\sim 0.1\,\%$ across the entire flux range. Similarly, the variation in the emission linewidth converges to $\sim 0.03\,\%$ for a Hilbert space dimension of $N=70$, at a representative flux of $\Phi=\Phi_{0}/4$, suggesting no significant impact on the overall system performance (further details can be found in the Supplementary Note 1). The Wigner functions defining the steady states of the reservoir for different values of the effective coupling $\mathrm{g_{eff}}$ are shown in Fig. 4(c). We observe the typical ”ring” shape that characterizes coherent photon sources, where the radius of the Wigner function is proportional to the average of the steady-state photon distribution [21, 37, 33]. After the reservoir has reached its steady state, the pumping process can be stopped and the flux $\mathrm{\Phi_{c}}$ set to zero. At this stage, to avoid further interactions between the photon source and the reservoir we can change the transition frequency of the photon source through the external flux $\mathrm{\Phi_{s}}$.
II.4 Multiple resonator generalization
@@
Another advantage offered by the proposed scheme is the ability to address multiple cavities by exploiting the tunability of the photon source. Under the same assumptions as the single-reservoir case, we can extend the analysis to an $\mathrm{N_{r}}$-reservoir system. The system is described by a Hamiltonian that includes additional terms related to the self-energy of different reservoirs, the coupling of each reservoir to the tunable coupler, and the second nearest neighbor cross-couplings between each resonator. In this discussion, we consider the simple case of a two-reservoir setup as a starting point, where the case of $\mathrm{N_{r}}$ reservoirs will be a direct generalization (calculations of the general $\mathrm{N_{r}}$ case are provided in Methods ”Schrieffer-Wolff transformation” ). We investigate the ability to individually address each reservoir through the external fluxes, threading the tunable coupler and assuming that the photon source can be tuned on resonance with both resonators separately. As in the single-reservoir case, a single coaxial line coupled to the tunable coupler can modulate the coupler frequency, enabling controlled emission to different target reservoirs. Assuming that both the coupler and the photon source behave as perfect two-level systems and neglecting second-order interactions, the Hamiltonian is written as:
$\mathrm{\hat{H}_{N}=\hat{H}_{s}+\hat{H}_{c}+\hat{H}_{r_{1}}+\hat{H}_{r_{2}}+\sum_{j=s,r_{1},r_{2}}\hat{H}_{jc}+\hat{H}_{r_{1}r_{2}}}\,.$ (8)
@@
$\mathrm{\frac{\hat{H}}{\hbar}}$ $=\mathrm{\frac{\tilde{\omega}_{s}\hat{\sigma}_{s}^{z}}{2}+\tilde{\omega}_{r_{1}}\hat{a}_{r_{1}}^{\dagger}\hat{a}_{r_{1}}+\tilde{\omega}_{r_{2}}\hat{a}_{r_{2}}^{\dagger}\hat{a}_{r_{2}}}$ (10)
$\mathrm{\phantom{=}+\sum_{j=r_{1},r_{2}}\tilde{g}_{jc}(\hat{\sigma}_{s}^{+}\hat{a}_{j}+\hat{\sigma}_{s}^{-}\hat{a}_{j}^{+})}$
$\mathrm{\phantom{=}+\tilde{g}_{r_{1}r_{2}}(\hat{a}_{r_{1}}\hat{a}^{\dagger}_{r_{2}}+\hat{a}_{r_{1}}^{\dagger}\hat{a}_{r_{2}})}\,.$
Figure 5: Multiple reservoirs case (a) Steady-state photon number in different cavities with resonance frequencies $\mathrm{\omega_{r_{1}}/2\pi=7\,GHz}$ and $\mathrm{\omega_{r_{2}}/2\pi=7.005\,GHz}$. The red and blue points on the two plots represent the average number of injected photons in each reservoir for specific choices of flux $\mathrm{\Phi_{c}}$ and frequency $\mathrm{\omega_{s}/2\pi}$ of the photon source. The absence of colored points in correspondence to the frequency of the other resonator in each colorplot shows the absence of cross-population phenomena. The circuital and simulation parameters are the same as the one considered in Figs. 2 and 4(b). (b) The Hausdorff distance between the effective coupling strength curve $\mathrm{g_{eff}(\Phi_{c})}$ of the reference reservoir and the second reservoir detuned by $\mathrm{\Delta}$ and coupled with a coupling capacity $\mathrm{\alpha}$ times that of the reference reservoir. The dark blue zone indicates a range of coupling capacities that maintain consistent effective coupling dependencies while altering the resonance frequency of the reservoir relative to the reference reservoir.
In this expression, the renormalizations of various frequency terms are identical to those observed in the single-reservoir case in Eq. 6, except for the new cross-coupling term $\mathrm{\tilde{g}_{r_{1}r_{2}}}$ between the two reservoirs. The second-order coupling term $\mathrm{g_{r_{1}r_{2}}}$ is summed with another term arising from the coupling between the reservoirs via the tunable coupler, leading to the following expression of the effective cross-coupling strength:
@@
It is, therefore, crucial to understand how much detuning is required between the different reservoirs at the fan-out of the system to avoid resonant interactions that lead to cross-population between the cavities. We consider a setup similar to the single-reservoir scheme analyzed in the previous section, with the addition of a second reservoir at a frequency $\mathrm{\omega_{r_{2}}=\omega_{r_{1}}+\Delta}$ and linked to the tunable coupler with the same capacitive coupling used for the target reservoir. We observe the system behavior for different values of $\mathrm{\Delta}$ while sweeping both the photon source and the coupler frequencies, simulating the effect of the external fluxes $\mathrm{\Phi_{s}}$ and $\Phi_{c}$ individually threading the two elements, respectively. Assuming that both reservoirs have the same linewidth as the reservoir considered in the previous single reservoir case $\mathrm{\kappa_{1}=\kappa_{2}=0.5\,MHz}$. To avoid cross-population between the two reservoirs the detuning $\mathrm{\Delta=\omega_{r_{1}}-\omega_{r_{2}}}$ must be much higher than the spectral linewidth of the two reservoirs $\mathrm{(\Delta\gg\kappa)}$. In this case, we chose a detuning of $\mathrm{\Delta/2\pi\geq 5\,MHz}$, that allows to individually address and control the steady-state population of each resonator with the proposed scheme.
In Fig. 5(a), we show the behavior of the system, where by setting the photon source on resonance with one of the cavities, it is possible to control their population without populating also the other resonator. For small detuning, the effective coupling between the source and the reservoir has approximately the same dependence on the flux, even for a fixed value of the capacitive coupling between the reservoir and the tunable coupler. When engineered systems are characterized by high detuning between the different reservoirs on the output, the respective capacitive links to the tunable coupler must be adequately designed. To optimize the device, the couplings must be chosen to preserve the shape of $\mathrm{g_{eff}(\Phi_{c})}$, because the effective coupling depends on the resonance frequency of the target reservoir $\mathrm{\omega_{t}}$. We thus need to take into account the variation of this frequency by changing the coupling between the central coupler and the reservoir under consideration. We use the same setup employed before, taking a reference reservoir, with a resonance frequency $\mathrm{\omega_{r_{1}}}$ and a capacity $\mathrm{C_{r_{1}c}}$ linking the reservoir to the central coupler. We consider a second reservoir with a detuned resonance frequency of $\mathrm{\omega_{r_{2}}=\omega_{r_{1}}+\Delta}$ and with a coupling capacitor $\mathrm{C_{r_{2}c}=\alpha C_{r_{1}c}}$, as depicted in Fig. 1(a). We assume that the self-capacity of the reservoir does not change with its resonance frequency. If there is a change it will result in an additional factor on the expression of the coupling strength that can be absorbed by the factor $\alpha$, allowing the same type of analysis. Through a sweep on $\mathrm{\Delta}$ and $\mathrm{\alpha}$, we study how the shape of $\mathrm{g_{eff}(\Phi_{c})}$ changes by analyzing the Hausdorff distance relative to the reference curve obtained for the reference reservoir. The results of this analysis are shown in Fig. 5(b), where we can clearly distinguish a blue region corresponding to the optimal values of the coupling capacity. For small values of the detuning $\mathrm{\Delta}$, the ideal coupling correction $\mathrm{\alpha}$ is near unity and, in general, follows the shape $\mathrm{\alpha=\sqrt{\omega_{r}/(\omega_{r}+\Delta)}}$.
@@
We have proposed a simple and versatile scheme for an on-demand coherent photon source on-chip. The tunability inherent to this scheme enables precise control over the average of the steady-state photon distribution generated in a target reservoir while also facilitating the addressing of different cavities capacitively linked to the coupler, each characterized by distinct resonance frequencies. Initially, we consider the case of a single resonator, where after identifying a set of parameters for a practical design we have analyzed the system performance for various fluxes applied across the coupler, demonstrating the capability of the proposed setup to control the photon distribution generated in a target reservoir. Subsequently, we extend our analysis to a setup comprising an additional reservoir, showcasing the capability to individually address different reservoirs at different frequencies by adequately designing the detunings and the couplings in the system. The device acts as a tunable and finely controllable cryogenic micromaser, which can be an important tool in many quantum architectures based on bosonic quantum systems, [53, 78, 79], like continuous variable quantum computing [80], quantum memories [81] and quantum metrology [82]. The fine degree of control over the emission also enables the implementation of this device to flip the state of superconducting qubits [40, 21] or to act as a master clock for microwave operations at the cryogenic stage [39]. Besides providing an external knob to control the emission from the photon source, the circuit defining the tunable coupler can be designed to enhance the emission linewidth [36, 38]. While linewidth optimization via coupling engineering falls beyond the scope of this work, it represents a promising direction for future studies, focused on engineering the coupling between the atom and the reservoir by incorporating non-linear elements [36, 38]. Moreover, the demultiplexing capability of the proposed scheme offers the advantage of decoupling multiple reservoirs from direct connections to room-temperature devices, leading to significant enhancements in terms of scalability. The scheme is also well-suited for 3D integrated systems, further expanding its domain of potential applications.
III Methods
III.1 Hamiltonian